OpenStudy (anonymous):

Given the hyperbola (x^2/4) – (y^2/9) = 1, find the equations for its asymptotes.

OpenStudy (anonymous):

A. x = (1/3)y, x = (–1/3)y B. y = (3/2)x, y = (–3/2)x C. y = (1/3)x, y = (–1/3)x D. x = (3/2)y, x = (–3/2)y

OpenStudy (anonymous):

Eq. of asymptotes (for right/left hyperbola): \[y= \pm \frac{b}{a}x\] so here, asymptotes : \[y= \pm \frac{3}{2}x\]